Weekly Experiment and Discussion - Part 1

  • Take 2 coins and flip "together" 50 times
  • Tally each set of flips and report the frequencies in the format shown below:
2nd coin Heads 2nd coin tails
1st coin Heads # of (H, H) # of (H, T)
1st coin Tails # of (T, H) # of (T, T)

(e.g., if on 12 occasions both coins were heads, there should be a 12 in the top left box)

Weekly Experiment and Discussion - Part 2

A) Given the results of your own experiment, can you conclude the results of each coin were independent? ( Show your supporting work)

B) Using the compiled results ( to be posted on Day #4), can you conclude the results of each coin were independent? ( Show your supporting work)


Note: For each, it is about interpreting the results versus any preconceived notions you may have about this exercise
Answer #1

Answer:

Let's take the case be:-

2nd coin Heads 2nd Coin Tail
1st coin Heads 13 17
1st coin Tails 8 12

The results of each coin are independent.

Define Event A = 1ST coin will show head

Event B = 2nd coin will show head

Events A & B are said to be independent iff P(A) * P(B) = P(A B)

Here P(A) = n(A)/n(S) = 13/(13+8) = 13/21

P(B) = n(B) / n(S) = 13/(13+ 17) = 13/ 30

P(A B) = n(A B) /n(S) = 13/ 50.

Here P(A) * P(B) = P(A B) = 0.2682 0.26 by probability approach, it is proved that eventa A & B are independent.

Using inferential Statistics we can prove that attributes A & B are independent by Chi Square teat of independence.

To test,

H0 : Attributes A & B are independent.

vs H1 : Attributes A & B are not independent.

It is 2X2 contingency table.

2nd coin head

2nd coin tail

Total

1st coin head

a = 13

b = 17

a+b = 30

1st coin tail

c = 8

d = 12

c+d = 20

total

a+c = 21

b+d = 29

N = 50

Under H0 , the test statistic is,

  

  

= (50* 400)/ 365400

= 200/3654

= 0.0547

From table the critical value for = 5% at 1 degrees of freedom is 3.841

Since calculated value of is less than that of tabulated value, we accept H0 .

CONCLUSION : The results of each coin were independent.